# Fair Document Valuation in LLM Summaries via Shapley Values

### v3

@@ 7.1 Comparison with Off-the-shelf Shapley Approximation Algorithms
    Kernel SHAP: Kernel SHAP is a model-agnostic approach to approximating Shapley values based on weighted least squares regression (40). Our implementation uses Python’s SHAP package, which we adapt specifically for our LLM-based summarization task by implementing a custom mapping function between subset compositions and their corresponding summarization scores. The method employs KernelExplainer with an identity link function and L1 regularization to enhance numerical stability. We tested increasing numbers of samples to evaluate the performance of the method under different computational budgets. Kernel SHAP has been widely used across domains, including NLP for transformer-based model interpretation (33), finance for credit risk analysis (16), healthcare for clinical decision support (37), and marketing for optimizing content engagement (34). Recent work has pointed out some limitations of Kernel SHAP, especially when interactions are strong (23; 53). [^13] Note that the issues pointed out by these papers relate to the approximation process used in the Kernel SHAP algorithm, and not the Shapley concept itself. Nevertheless, given that it is one of the most widely used Shapley approximations used in practice, we include it in our evaluation for benchmarking purposes.
@@ 8 Conclusion
Another promising direction is to consider how document value attribution systems can be combined with reputation mechanisms. While a large stream of theoretical and empirical research has established the importance of reputation mechanisms in sustaining two-sided platforms (56; 12; 70), it is unclear how these systems can survive without direct feedback on contributor’s performance. Thus, extending the document valuation framework to augment/replace existing reputation systems can be an important next step.
@@ B.2 Empirical Validation of Lipschitz Continuity Assumption
The scatter plot demonstrates that marginal contribution differences are linearly bounded by embedding distances, providing strong empirical support for the Lipschitz continuity assumption. The data reveals a clear pattern: document pairs with small embedding distances (high similarity) consistently exhibit small differences in their marginal contributions across coalitions. Conversely, document pairs with large embedding distances show a wider range of marginal contribution differences, from small to large values.
This asymmetric relationship aligns with our theoretical expectations and practical intuition. When two documents have similar embeddings, they naturally contribute similarly to any coalition, resulting in consistently small marginal contribution differences. However, when documents have dissimilar embeddings, their marginal contribution differences can vary significantly depending on the specific aspects they emphasize. For instance, two very different documents describing distinct aspects of a product (e.g., technical specifications versus user experience) may still provide comparable value to certain coalitions, leading to small marginal contribution differences despite large embedding distances.
The observed upper bound behavior-where embedding distance provides a ceiling for marginal contribution differences-validates the Lipschitz continuity assumption while capturing the nuanced relationship between semantic similarity and functional contribution in our evaluation framework. This empirical evidence supports the use of embedding-based distances as meaningful constraints in Shapley value computations.

### v4

@@ 7.1 Comparison with Off-the-shelf Shapley Approximation Algorithms
    Kernel SHAP: Kernel SHAP is a model-agnostic approach to approximating Shapley values based on weighted least squares regression (47). Our implementation uses Python’s SHAP package, which we adapt specifically for our LLM-based summarization task by implementing a custom mapping function between subset compositions and their corresponding summarization scores. The method employs KernelExplainer with an identity link function and L1 regularization to enhance numerical stability. We tested increasing numbers of samples to evaluate the performance of the method under different computational budgets. Kernel SHAP has been widely used across domains, including NLP for transformer-based model interpretation (40), finance for credit risk analysis (22), healthcare for clinical decision support (44), and marketing for optimizing content engagement (41). Recent work has pointed out some limitations of Kernel SHAP, especially when interactions are strong (29; 61). [^14] Note that the issues pointed out by these papers relate to the approximation process used in the Kernel SHAP algorithm, and not the Shapley concept itself. Nevertheless, given that it is one of the most widely used Shapley approximations used in practice, we include it in our evaluation for benchmarking purposes.
@@ 8 Conclusion
Another promising direction is to consider how document value attribution systems can be combined with reputation mechanisms. While a large stream of theoretical and empirical research has established the importance of reputation mechanisms in sustaining two-sided platforms (64; 17; 80), it is unclear how these systems can survive without direct feedback on contributor’s performance. Thus, extending the document valuation framework to augment/replace existing reputation systems can be an important next step.
@@ B.2 Empirical Validation of Lipschitz Continuity Assumption
The scatter plot demonstrates that marginal contribution differences are linearly bounded by embedding distances, providing strong empirical support for the Lipschitz continuity assumption. The data reveals a clear pattern: document pairs with small embedding distances (high similarity) consistently exhibit small differences in their marginal contributions across coalitions. Conversely, document pairs with large embedding distances show a wider range of marginal contribution differences, from small to large values. This asymmetric relationship aligns with our theoretical expectations and practical intuition. When two documents have similar embeddings, they naturally contribute similarly to any coalition, resulting in consistently small marginal contribution differences. However, when documents have dissimilar embeddings, their marginal contribution differences can vary significantly depending on the specific aspects they emphasize. For instance, two very different documents describing distinct aspects of a product (e.g., technical specifications versus user experience) may still provide comparable value to certain coalitions, leading to small marginal contribution differences despite large embedding distances.
Next, we further quantify this relationship. Recall that the y-axis represents the difference in marginal contributions, defined as:
$\Delta_{i,j}(S)=\big|\big(v(S\cup\{i\})-v(S)\big)-\big(v(S\cup\{j\})-v(S)\big)\big|.$
Since our evaluation prompt restricts the performance score $v(\cdot)$ to the integer range $[0,10]$, $\Delta_{i,j}(S)$ is naturally strictly bounded by 10. Consequently, a global linear bound with a slope of $L=10$ is somewhat trivial (given that the embedding distance is normalized to $[0,1]$) and practically a very loose bound. However, the validity of Assumption 1 hinges on the local behavior of this difference as the embedding distance approaches zero. If the ratio of differences to distances were unbounded, we would observe data points filling the upper-left region of the plot (i.e., small distance but large difference).
Instead, Figure A6 shows a dense concentration of points near the origin. To quantify this, we examine the data points within the region of high semantic similarity (specifically, embedding distance $d(e_{i},e_{j})\in[0,0.4]$). In this range, we find that more than 95% of the data points satisfy the condition:
$\Delta_{i,j}(S)\leq 2.5\times d(e_{i},e_{j}).$
This indicates that for the vast majority of cases, the effective Lipschitz constant is approximately $L\approx 2.5$, which is significantly tighter than the trivial global bound of 10. While a small fraction of outliers exists due to the inherent stochasticity of LLM scoring (e.g., randomness in generation or evaluation).